Small Sample Mean Problem. The maker of potato chips uses an automated packaging machine to pack its 20-ounce bag of chips. At the end of a shift 15 bags are selected at random and tested to see if the equipment needs to be readjusted. After one shift, a sample of 15 bags yielded the following data. mean = 19.45 s = .64 n = 15. A 90% confidence interval would have what as the lower and upper bounds?

0?ui=2&ik=171117cc57&attid=0.1&permmsgid

0?ui=2&ik=171117cc57&attid=0.1&permmsgid

0?ui=2&ik=171117cc57&attid=0.1&permmsgid

0?ui=2&ik=171117cc57&attid=0.1&permmsgid
Answer #1

Sample mean: Sample size: Sample standard deviation: x = n= s= n-1= a= a/2 = 19.45 given 15 given 0.64 given 14 0.1 given 90%